Using imlicit differentiation, which one of the following is the derivative of

"cos(x-2y) = sin(2x + y)"


1
Expert's answer
2021-06-07T17:25:32-0400

"cos(x−2y)=sin(2x+y)\\
\text{differentiate with respect to x.}\\
\implies -sin(x−2y)(1-2y')=cos(2x+y)(2+y')\\
\implies
-sin(x−2y)+2sin(x−2y)y'-2cos(2x+y)-
cos(2x+y)y'=0\\
\implies
y'=\frac{sin(x−2y)+2cos(2x+y)}{2sin(x−2y)-cos(2x+y)}"